Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#248 closed flight processing (fixed)

HY05/06, flight day 152b/2009, Pitlocry

Reported by: jatt Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by emca)

Data location: ~arsf/arsf_data/2009/flight_data/uk/HY05_06-2009_152b_Pitlocry

Data arrived from ARSF via SATA disk M on 9th June 2009.

Scientific objective: N/K

Priority: α4L

PI: M Cutler

Sensors:

  • Eagle (11/11/2009)
  • Hawk (11/11/2009)
  • Leica LIDAR (25/05/2010)
  • Digital Photography (09/02/2010)

Change History (33)

comment:1 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:2 Changed 10 years ago by knpa

Applanix Processing Completed

comment:3 Changed 10 years ago by knpa

No sync info for E+H line 6

comment:4 Changed 10 years ago by knpa

Semi-regular line drops in Hawk, around two per line flightline of approx 8 scan lines thickness.

comment:5 Changed 10 years ago by knpa

Large line drop on flight-line e1, approx 40 scan lines in thickness.

comment:6 Changed 10 years ago by knpa

E+H Processing completed, awaiting vectors.

comment:7 Changed 10 years ago by knpa

Contacted PI and they are happy to wait for vectors.

comment:8 Changed 10 years ago by mark1

Vectors have arrived.

comment:9 Changed 10 years ago by knpa

Vector check shows generally good positioning, however appears to be a horizontal sift of around 4m across indicated by roads on multiple flightlines.

comment:10 Changed 10 years ago by knpa

the above note seems to apply to eagle only

comment:11 Changed 10 years ago by mark1

  • Owner changed from knpa to mark1
  • Status changed from assigned to new

comment:12 Changed 10 years ago by mark1

Delivery checking.

Minor edits made to readme. Example commands tested (without dem) and run ok. All level 1 files checked and look good.

comment:13 Changed 10 years ago by knpa

  • Summary changed from HY05/06, flight day 152b/2009, Pitlocry to added sct timings

derived timing offsets for this project:

SCT OFFSET


Eagle Hawk

1 0.00 0.00
2 0.06 0.07
3 0.03 0.04
4 0.08 0.01
5 0.00 0.05
6 0.70 1.08

comment:14 Changed 10 years ago by knpa

  • Summary changed from added sct timings to HY05/06, flight day 152b/2009, Pitlocry

comment:15 Changed 10 years ago by knpa

E H
10.00 0.00
20.06 0.07
30.03 0.04
40.08 0.01
5 0.00 0.05
60.70 1.08

comment:16 Changed 10 years ago by mark1

  • Owner changed from mark1 to knpa

data rsynced over to ~arsf/arsf_data/2009/flight_data/uk/HY05_06-2009_152b_Pitlocry

comment:17 Changed 10 years ago by knpa

HYPERSPECTRAL DELIVERY TODAY

comment:18 Changed 10 years ago by knpa

  • Owner knpa deleted

comment:19 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:20 Changed 10 years ago by mark1

Have edited the 090601-235947ImageEvents1.csv file (lines 28 and 40) to put in an approximate "GPS time" (calculated from 090601-235947SensorStat1.csv file) to allow the image to be tagged with data. The times from 090601-235947SensorStat1.csv appear to be accurate (compared to the GPS time) to within 0.01s.

comment:21 Changed 10 years ago by mark1

Photographs dispatched to PI on 9th February 2010

comment:22 Changed 9 years ago by mark1

  • Description modified (diff)

comment:23 Changed 9 years ago by emca

  • Owner changed from knpa to emca

Starting LIDAR processing

comment:24 Changed 9 years ago by emca

LIDAR processing complete. Ready for delivery check at
~airborne/workspace/HY05_06-2009_152b_Pitlocry/delivery/20100519/HY05_06

comment:25 Changed 9 years ago by mark1

LIDAR delivery checked and all looks good.
Ready to deliver.

comment:26 Changed 9 years ago by emca

LIDAR data sent to PI 25/05/2010

comment:27 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

LIDAR data rsynced to:
~arsf/arsf_data/2009/flight_data/uk/HY05_06-2009_152b_Pitlocry

Photography and Hyperspectral already present.

Ready for archiving.

comment:28 Changed 9 years ago by knpa

  • Owner changed from emca to knpa

Beginning archiving.

comment:29 Changed 9 years ago by knpa

NEODC have downloaded and backed up.

Moved to archive disk.

comment:30 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:31 Changed 6 years ago by stgo

SCTs from config file

Flightline Eagle Hawk
1 -0.02 -0.02
2 -0.01 -0.02
3 -0.03 0.00
4 -0.03 -0.05
5 -0.09 0.92
6 -0.69 1.00

comment:32 Changed 6 years ago by stgo

PPP result

lat: 57 00 24.25265
long: -3 23 44.16777
ell. height: 401.077

comment:33 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.